1/1

Михаил Юрьевич Лермонтов, Михаил Юрьевич Лермонтов,

Михаил Юрьевич Лермонтов,

Парус  Das Segel  Sail
tradukita de B. Denken  tradukita de A. Z. Foreman

Белеет парус одинокий В тумане моря

голубом!.. Что ищет он в стране

далекой? Что кинул он в краю

родном?..

Verloren glänzt in blauen Weiten Ein Segel weiß in dunst’gem

Schein. Was sucht es nur in fernen Brei-

ten? Was ließ es Hort und Heim al-

lein?

A sail drifts white and on its own Amid the light blue ocean haze. What does it seek in distant coun-

try? What made it leave its native

bays?

Играют волны, ветер

свищет, И мачта гнётся и

скрипит, Увы! – он счастия не

ищет, И не от счастия бежит! – Es spiel’n die Wogen, Winde pfei-

fen, Es jammert, ächzt und knirscht

der Mast – Doch, ach! nach Glück will es

nicht greifen Noch ihm entflieh’n in blinder

Hast.

Great billows play. High winds are

whistling Down at the bending, creaking

mast Oh! This one seeks no happy en-

ding And does not flee a happy past.

Под ним струя светлей

лазуры, Над ним луч солнца

золотой: – А он мятежный, просит

бури, Как будто в бурях есть

покой.

Tief unten schäumt lichtblau die

Flut, Hoch oben gleißen Sonnes

Brände, Doch sucht es mehr nach Stur-

mes Wut, Als ob’s im Sturme — Ruhe

fände.

Beneath, a brighter stream than

azure. Above, the golden sunray flows Yet this one, restive...quests for

tempests As if in tempests were repose.

Verkinto de tiu ĉi Rusa poemo estas Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-27).

Arg-545-1093 (2010-06-20 16:02:28)

Traduko de la Rusa poemo “Парус” de Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-27) en la Germanan de B. Denken.

Arg-545-1094 (2010-06-20 16:15:25)

Tiun ĉi poem-germanigon mi, Manfred Retzlaff, trovis en la retejo http: // www. medeasy. de/ common/ ?p= 1349 .

Traduko de la Rusa poemo “Парус” de Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-27) en la Anglan de A. Z. Foreman.

Arg-545-2269 (2015-01-28 19:03:20)

I, Manfred Retzlaff, found this poem-translation in http: // poemsintranslation. blogspot. e/ 2009/ 10/ lermontov-sail-from-russian. html .

http://www.poezio.net/version?poem-id=545&version-id=1093,1094,2269